# Towards a Sharp Analysis of Offline Policy Learning for $f$-Divergence-Regularized Contextual Bandits

### v2

@@ D.1 Algorithms for KL-regularized Contextual Dueling Bandits
2:  Let ${\widehat{g}}(s,a)={\bar{g}}(s,a)-\Gamma_{n}(s,a)$, where $\Gamma_{n}(s,a)$ is the bonus term in D.1
2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{g}}(s,a)\big{)}$
@@
$\Gamma_{n}(s,a)=\beta\sqrt{D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})},$ (D.1)
@@
$\beta^{2}=128\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)/3n+18\epsilon_{c}=\widetilde{O}(n^{-1})$ (D.2)
and then subtract it from the MLE ${\bar{g}}$ to obtain a pessimistic estimator ${\widehat{g}}$. KL-PCB then output the policy $\widehat{\pi}$, maximizing the estimated objective
@@
Theorem D.1.
Under Assumption 4.3, if we set $\Gamma_{n}$ according to D.1, then for sufficiently small $\epsilon\in(0,1)$, with probability at least $1-\delta$, $n=\widetilde{O}\big{(}\eta D_{\pi^{*}}^{2}\epsilon^{-1}\big{)}$ is sufficient to guarantee the output policy $\widehat{\pi}$ of Algorithm 3 to be $\epsilon$-optimal.
Remark D.2.
Zhao et al. (2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption 4.2. Comparing to Zhao et al. (2024), KL-PCDB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption 4.3, which is weaker than Assumption 4.2.
@@
Theorem D.3.
For any sufficiently small $\epsilon\in(0,1),\eta>0$, $21\leq C^{*}\leq\exp(\eta/2)/$ 2, and any algorithm Alg, there is a KL-regularized contextual dueling bandit instance with single-policy concentrability $D^{2}_{\pi^{*}}\leq 2C^{*}$ such that Alg requires at least $\Omega\big{(}\min\{{\eta C^{*}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})}/{\epsilon},{\log\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})}(C^{*})^{2}/{\epsilon^{2}}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
Remark D.4.
Theorem D.3 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual dueling bandits requires at least $\Omega(\eta D^{2}_{\pi^{*}}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\epsilon^{-1})$ samples to output an $\epsilon$-optimal policy, which matches the sample complexity upper bound in Theorem D.1, indicating that KL-PCB is nearly optimal.
D.2 Algorithm and Results for $f$-divergence Regularized Contextual Dueling Bandit
@@
2:  $\widehat{\pi}(a|s)$
In this subsection, we present an offline learning algorithm for $f$-divergence regularized contextual dueling bandit, $f$-CDB, in Algorithm 4. $f$-CDB first leverages maximum likelihood estimator to find a function ${\bar{g}}\in{\mathcal{G}}$ that minimizes its risk on the offline dataset. Then the algorithm constructs the output policy $\widehat{\pi}$ that maximizes the f-divergence regularized objective induced by ${\bar{g}}$. Similar to Algorithm 2, we do not require any pessimism in $f$-CDB. The following theorem provides an upper bound of Algorithm 4.
Theorem D.5.
For any sufficiently small $\epsilon\in(0,1)$, and $\eta,\alpha>0$, with probability at least $1-\delta$, $n=\widetilde{O}\big{(}\alpha^{-1}\eta\log{\mathcal{N}}(\epsilon)\epsilon^{-1}\big{)}$ is sufficient to guarantee that the output policy $\widehat{\pi}$ of Algorithm 4 is $\epsilon$-optimal.
@@
Remark D.7.
Theorem D.6 indicates that, when $\epsilon$ is sufficiently small, to produce an $\epsilon$-optimal policy, any algorithm for offline $f$-regularized contextual bandits with strongly convex $f$ requires at least $\widetilde{\Omega}(\alpha^{-1}\eta\epsilon^{-1})$ samples. This lower bound matches the sample-complexity upper bound in Theorem D.5, indicating that Algorithm 4 is nearly optimal.
Appendix E Missing Proof from Appendix D
E.1 Proof of Theorem D.1
@@ Appendix E Missing Proof from Appendix D
$\mathcal{E}(\delta)\coloneqq\Big{\{}\exists\ b:{\mathcal{S}}\to[-1,1],\forall(s,a)\in{\mathcal{S}}\times\mathcal{A},\big{|}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{|}\leq\Gamma_{n}(s,a)\Big{\}}.$ (E.1)
Here, $\Gamma_{n}$ is defined in (D.1). We abuse the notation and define $b(\cdot)$ as
$b=\mathop{\mathrm{argmin}}_{\mathcal{B}}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Phi_{b}(s,a)-\Gamma_{n}(s,a),$ (E.2)
@@
From Lemma F.1, we have that with probability at least $1-\delta$, it holds that
$\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{g}}(s^{\prime},a^{\prime})-{g^{*}}(s^{\prime},a^{\prime})\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$ (E.3)
@@
$D_{\mathcal{G}}^{2}((s,a),\pi^{\mathsf{ref}}))\cdot\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a\sim\pi^{\mathsf{ref}}(\cdot|s)}\big{[}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{]}\geq\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}.$ (E.4)
Substituting (E.3) into (E.4), we have
$\inf_{b}\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}$ (E.5)
@@
$\quad\leq D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}\big{]}$ (E.6)
$\quad\leq D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)},$ (E.7)
where the first inequality holds due to the definition of $D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})$ and the last inequality holds due to Lemma F.1. ∎
@@
where $b(\cdot)$ is defined in (E.2). We still have the monotonicity of the function $G(\gamma)$, which is characterized by the following lemma.
Lemma E.2.
@@
Proposition E.3.
Let $b:{\mathcal{S}}\to\mathbb{R}$ be some bias function, then for all $g\in{\mathcal{G}}$ we have $J(\pi_{g})=J(\pi_{g-b})$, where $(g-b)(s,a)=g(s,a)-b(s)$.
@@
Now we are ready to prove Theorem D.1.
Proof of Theorem D.1.
We proceed the proof under the event $\mathcal{E}(\delta)$. By Proposition E.3, we know that
$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{g}}})$
@@
$=\eta G(\gamma),$ (E.8)
where the inequality holds due to Lemma 2.16. Under event $\mathcal{E}(\delta)$, we know that ${\widehat{g}}(s,a)-b(s)\leq{{g^{*}}}(s,a)$. Together with Lemma E.2, we obtain $G(\gamma)\leq G(0)$. Therefore, we know that
$J(\pi^{*})-J(\widehat{\pi})$ $\leq G(0)$
@@
$\leq 4\eta\mathbb{E}_{\rho\times\pi^{*}}\big{[}\Gamma_{n}^{2}(s,a)\big{]}$
$=4\eta\beta^{2}\mathbb{E}_{\rho\times\pi^{*}}\big{[}D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})\big{]}$
$=\widetilde{O}\big{(}\eta D_{\pi^{*}}^{2}\log{{\mathcal{G}}}(\epsilon_{c})n^{-1}\big{)},$ (E.9)
where the inequality holds due to the definition of $\mathcal{E}(\delta)$. Plugging (E.9) into (E.8), we know that $J(\pi^{*})-J(\widehat{\pi})$ has upper bound $\widetilde{O}(D_{\pi^{*}}^{2}n^{-1})$. By Lemma E.1, event $\mathcal{E}$ with probability at least $1-\delta$, which concludes the proof. ∎
E.2 Proof of Theorem D.3
Proof of Theorem D.3.
@@
$\mathrm{CDB}\coloneqq\{({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta):r\in\mathcal{G},\rho\in\Delta({\mathcal{S}}),\pi^{\mathsf{ref}}\in\Delta(\mathcal{A}|{\mathcal{S}})\}.$ (E.10)
Fixing any $S\geq 1$, $\eta>4\log 2$ and $C^{*}\in(2,\exp(\eta/4)]$, we aim to prove that, for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n\geq 16SC^{*}$, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CDB}$ with single-policy concentrability $C^{\pi^{*}}\leq 2C^{*}$, regularization coefficient $\eta$, $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
$\inf_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})\gtrsim\min\{\eta SC^{*}n^{-1},(SC^{*})^{1/2}n^{-1/2}\}.$ (E.11)
@@
[…]
@@
$\mathrm{SubOpt}_{f\mathrm{div}}(\pi;r)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{f\mathrm{div}}\big{(}\pi(\cdot|s);r(s,\cdot)\big{)}.$ (E.12)
@@
Then, following the derivation of B.10 and B.11, we know that $\max_{\tau\in\{\pm 1\}^{S}}C^{\pi_{\tau}^{*}}\leq C^{*}$, and $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
$\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)+\mathrm{SubOpt}_{s}(\widehat{\pi};\tau^{\prime})\geq\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}.$ (E.13)
Let $P_{r}$ be the distribution of $(s,a^{1},a^{2},y)$ for $s\sim\rho$, $a^{1},a^{2}\overset{\scriptsize\mathrm{i.i.d.}}{\sim}\pi^{\mathsf{ref}}(\cdot|s)$ and $y\sim\mathsf{Bern}(\sigma(r(s,a^{1})-r(s,a^{2})))$. Now we set $\delta=\sqrt{S/n}$ and conclude that for $\tau\sim\tau^{\prime}$ with $\tau_{s}=-\tau_{s}^{\prime}$,
@@
$\leq 36S^{-1}C^{-1}\delta^{2},$ (E.14)
where the second and third inequality hold due to $\alpha,\delta\leq 1/2$, and last inequality follows from $\exp(x)-\exp(-x)\leq 3x$ for $x\in[0,1/2]$. Now we set $\delta=\sqrt{SC/n}\leq 1/4$. We substitute E.13 into Assouad’s Lemma (Lemma F.3) and obtain that
@@
where the $1/S$ comes from the denominator of (E.12) and the second inequality follows from E.14. ∎
E.3 Proof of Theorem D.5
Proof of Theorem D.5.
The proof is similar to the proof of Theorem 3.2. Recall that $b(\cdot)$ defined in (E.2), we know that
@@
$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{s\sim\rho}\Big{[}\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(s,a)]-\mathbb{E}_{a\sim\widehat{\pi}}[{g^{*}}(s,a)]-\eta^{-1}\big{[}D_{f}(\pi^{*}\|\pi^{\mathsf{ref}})-D_{f}(\widehat{\pi}\|\pi^{\mathsf{ref}})\big{]}\Big{]}$
@@
where the first inequality holds due to Lemma C.4 and last inequality holds due to equation (E.3). Setting $\epsilon_{c}=O(n^{-1})$ completes the proof. ∎
E.4 Proof of Theorem D.6
Proof of Theorem D.6.
We still consider the contextual dueling bandit instance class defined in (E.10). We show that given any positive $\alpha,\eta$, for any $n\geq S\cdot\max\{16,\eta^{2}\alpha^{-2}\}$, there exists $f:\mathbb{R}\to\mathbb{R}$ such that $f$ is $\alpha$-strongly convex, $\log|\mathcal{G}|=\Theta(S)$ and
$\inf_{\widehat{\pi}\in\widehat{\Pi}(\mathcal{D})}\sup_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{f\mathrm{div}}(\widehat{\pi};\mathrm{inst})\gtrsim\frac{\eta S}{\alpha n},$ (E.15)
@@
[…]
$\mathrm{SubOpt}_{f\mathrm{div}}(\pi;r)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{f\mathrm{div}}\big{(}\pi(\cdot|s);r(s,\cdot)\big{)}.$ (E.16)
We still consider the reward function class $\mathcal{G}$ indexed by {±$\{\pm 1\}^{{\mathcal{S}}}$. For all $\tau\in\{\pm 1\}^{{\mathcal{S}}}$ the reward instance “shaped” by $\tau$ is
$r_{\tau}(s,a)=\frac{1}{2}+a\tau_{s}\cdot\sqrt{\frac{S}{n}},$ (E.17)
where $a\tau_{s}=\pm 1$ because $a\in\mathcal{A}=\{\pm 1\}$. We thereby refer $\tau\sim\tau^{\prime}$ to any pair in {±$\{\pm 1\}^{{\mathcal{S}}}$ that differs only in one coordinate. $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$, if $\tau\sim\tau^{\prime}$, then suppose $\tau_{s}=-\tau_{s}^{\prime}$, we have
$\mathrm{SubOpt}_{f\mathrm{div}}(\pi(\cdot|s);r_{\tau}(s,\cdot))+\mathrm{SubOpt}_{f\mathrm{div}}(\pi(\cdot|s);r_{\tau^{\prime}}(s,\cdot))\geq\frac{\eta S}{\alpha n},$ (E.18)
[…]

### v3

@@ F.2.1 Algorithms for KL-regularized contextual dueling bandits
2: Let ${\widehat{g}}(s,a)={\bar{g}}(s,a)-\Gamma_{n}(s,a)$, where $\Gamma_{n}(s,a)$ is the bonus term in F.1
[…]
@@
$\Gamma_{n}(s,a)=\beta\sqrt{D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})},$ (F.1)
@@
$\beta^{2}=128\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)/3n+18\epsilon_{c}=\widetilde{O}(n^{-1})$ (F.2)
and then subtract it from the MLE $\bar{g}$} to obtain a pessimistic estimator $^{\widehat{g}}$. KL-PCB then output the policy $\widehat{\pi}$, maximizing the estimated objective
@@
Theorem F.4.
Under Assumption F.3, if we set $\Gamma_{n}$ according to F.1, then for sufficiently small $\epsilon\in(0,1)$, with probability at least $1-\delta$, $n=\widetilde{O}\big(\eta(D_{\pi^{*}}^{2}\wedge C^{\pi^{*}})\epsilon^{-1}\big)$ is sufficient to guarantee the output policy $\widehat{\pi}$ of Algorithm 3 to be $\epsilon$-optimal.
Remark F.5.
Zhao et al. (2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption F.2. Comparing to Zhao et al. (2024), KL-PCDB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption F.3, which is weaker than Assumption F.2.
@@
Theorem F.6.
For any sufficiently small $\epsilon\in(0,1),\eta>0$, $1\leq C^{*}\leq\exp(\eta/2)/2$, and any algorithm Alg, there is a KL-regularized contextual dueling bandit instance with single-policy concentrability $C^{\pi^{*}}\leq C^{*}$ such that Alg requires at least $\Omega\big(\min\{{\eta C^{*}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})}/{\epsilon},{\log\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})}(C^{*})^{2}/{\epsilon^{2}}\}\big)$ samples to return an $\epsilon$-optimal policy.
Remark F.7.
Theorem F.6 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual dueling bandits requires at least $\Omega(\eta C^{\pi^{*}}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\epsilon^{-1})$ samples to output an $\epsilon$-optimal policy, which matches the sample complexity upper bound in Theorem F.4, indicating that KL-PCB is nearly optimal.
F.2.2 Algorithm and Results for $f$-divergence Regularized CDBs
@@
[…]
We present an offline learning algorithm for $f$-divergence regularized contextual dueling bandit, $f$-CDB, in Algorithm 4. $f$-CDB first leverages maximum likelihood estimator to find a function ${\bar{g}}\in{\mathcal{G}}$ that minimizes its risk on the offline dataset. Then the algorithm constructs the output policy $\widehat{\pi}$ that maximizes the f-divergence regularized objective induced by $\bar{g}$}. Similar to Algorithm 2, we do not require any pessimism in $f$-CDB. The following theorem provides an upper bound of Algorithm 4.
Theorem F.8.
For any sufficiently small $\epsilon\in(0,1)$, and $\eta,\alpha>0$, with probability at least $1-\delta$, $n=\widetilde{O}(\alpha^{-1}\eta\log{\mathcal{N}}(\epsilon)\epsilon^{-1})$ is sufficient to guarantee that the output policy $\widehat{\pi}$ of Algorithm 4 is $\epsilon$-optimal.
@@
Remark F.10.
Theorem F.9 indicates that, when $\epsilon$ is sufficiently small, to produce an $\epsilon$-optimal policy, any algorithm for offline $f$-regularized contextual bandits with strongly convex $f$ requires at least $\widetilde{\Omega}(\alpha^{-1}\eta\epsilon^{-1})$ samples. This lower bound matches the sample-complexity upper bound in Theorem F.8, indicating that Algorithm 4 is nearly optimal.
Appendix G Missing proof from Appendix F
G.1 Proof of Theorem F.4
@@ Appendix G Missing proof from Appendix F
$\mathcal{E}(\delta)\coloneqq\Big\{\exists\ b:{\mathcal{S}}\to[-1,1],\forall(s,a)\in{\mathcal{S}}\times\mathcal{A},\big|{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big|\leq\Gamma_{n}(s,a)\Big\}.$ (G.1)
Here, $\Gamma_{n}$ is defined in (F.1). We abuse the notation and define $b(\cdot)$ as
$b=\mathop{\mathrm{argmin}}_{\mathcal{B}}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Phi_{b}(s,a)-\Gamma_{n}(s,a),$ (G.2)
@@
From Lemma H.1, we have that with probability at least $1-\delta$, it holds that
$\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big[{\bar{g}}(s^{\prime},a^{\prime})-{g^{*}}(s^{\prime},a^{\prime})\big]\leq O\bigg(\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg).$ (G.3)
@@
$D_{\mathcal{G}}^{2}((s,a),\pi^{\mathsf{ref}}))\cdot\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a\sim\pi^{\mathsf{ref}}(\cdot|s)}\big[{\bar{g}}(s,a)-{g^{*}}(s,a)\big]\geq\big({\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big)^{2}.$ (G.4)
Substituting (G.3) into (G.4), we have
$\inf_{b}\big({\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big)^{2}$ (G.5)
@@
$\quad\leq D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big[\big({\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big)^{2}\big]$ (G.6)
$\quad\leq D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})O\bigg(\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg),$ (G.7)
where the first inequality holds due to the definition of $D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})$ and the last inequality holds due to Lemma H.1. ∎
@@
where $b(\cdot)$ is defined in (G.2). We still have the monotonicity of the function $G(\gamma)$, which is characterized by the following lemma.
Lemma G.2.
@@
Proposition G.3.
Let $b:{\mathcal{S}}\to\mathbb{R}$ be some bias function, then for all $g\in{\mathcal{G}}$ we have $J(\pi_{g})=J(\pi_{g-b})$, where (gb)(s,a)=g(s,a)b(s)(g-b)(s,a)=g(s,a)-b(s).
@@
Now we are ready to prove Theorem F.4.
Proof of Theorem F.4.
We proceed the proof under the event $\mathcal{E}(\delta)$. By Proposition G.3, we know that
[…]
@@
$=\eta G(\gamma),$ (G.8)
where the inequality holds due to Lemma 2.14. Under event $\mathcal{E}(\delta)$, we know that ${\widehat{g}}(s,a)-b(s)\leq{{g^{*}}}(s,a)$. Together with Lemma G.2, we obtain $G(\gamma)\leq G(0)$. Therefore, we know that
[…]
@@
$\leq 4\eta\Big(\mathbb{E}_{\rho\times\pi^{*}}\big[\Gamma_{n}^{2}(s,a)\big]\wedge C^{\pi^{*}}\mathbb{E}_{\rho\times\pi^{*}}\big[\big({\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big)^{2}\big]\Big)$
$=4\eta\Big(\beta^{2}\mathbb{E}_{\rho\times\pi^{*}}\big[D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})\big]\wedge C^{\pi^{*}}\mathbb{E}_{\rho\times\pi^{*}}\big[\big({\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big)^{2}\big]\Big)$
$=\widetilde{O}\big(\eta D_{\pi^{*}}^{2}\log{{\mathcal{G}}}(\epsilon_{c})n^{-1}\big),$ (G.10)
[…]
G.2 Proof of Theorem F.6
Proof of Theorem F.6.
@@
$\mathrm{CDB}\coloneqq\{({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta):r\in\mathcal{G},\rho\in\Delta({\mathcal{S}}),\pi^{\mathsf{ref}}\in\Delta(\mathcal{A}|{\mathcal{S}})\}.$ (G.11)
Fixing any $S\geq 1$, $\eta>4\log 2$ and $C^{*}\in(2,\exp(\eta/4)]$, we aim to prove that, for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n\geq 16SC^{*}$, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CDB}$ with single-policy concentrability $C^{\pi^{*}}\leq C^{*}$, regularization coefficient $\eta$, $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
$\inf_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})\gtrsim\min\{\eta SC^{*}n^{-1},(SC^{*})^{1/2}n^{-1/2}\}.$ (G.12)
@@
We construct the same reward function class as in the proof of Theorem 2.11. In particular, we set ${\mathcal{S}}=[S]$, $\mathcal{A}=\{\pm 1\}$, $\rho=\mathsf{Unif}({\mathcal{S}})$, and the reference policy to be
@@
$\mathrm{SubOpt}_{f\mathrm{div}}(\pi;r)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{f\mathrm{div}}\big(\pi(\cdot|s);r(s,\cdot)\big).$ (G.13)
@@
Then, following the derivation of D.12 and D.13, we know that $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
$\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)+\mathrm{SubOpt}_{s}(\widehat{\pi};\tau^{\prime})\geq\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}.$ (G.14)
Let $P_{r}$ be the distribution of $(s,a^{1},a^{2},y)$ for $s\sim\rho$, $a^{1},a^{2}\overset{\mathrm{i.i.d.}}{\sim}\pi^{\mathsf{ref}}(\cdot|s)$ and $y\sim\mathsf{Bern}(\sigma(r(s,a^{1})-r(s,a^{2})))$. Now we set $\delta=\sqrt{S/n}$ and conclude that for $\tau\sim\tau^{\prime}$ with $\tau_{s}=-\tau_{s}^{\prime}$,
@@
$\leq 36S^{-1}C^{-1}\delta^{2},$ (G.15)
where the second and third inequality hold due to $\alpha,\delta\leq 1/2$, and last inequality follows from $\exp(x)-\exp(-x)\leq 3x$ for $x\in[0,1/2]$. Now we set $\delta=\sqrt{SC/n}\leq 1/4$. We substitute G.14 into Assouad’s Lemma (Lemma H.3) and obtain that
@@
where the $1/S$ comes from the denominator of (G.13) and the second inequality follows from G.15. ∎
G.3 Proof of Theorem F.8
Proof of Theorem F.8.
The proof is similar to the proof of Theorem 3.2. Recall that $b(\cdot)$ defined in (G.2), we know that
@@
[…]
@@
where the first inequality holds due to Lemma E.4 and last inequality holds due to equation (G.3). Setting $\epsilon_{c}=O(n^{-1})$ completes the proof. ∎
G.4 Proof of Theorem F.9
Proof of Theorem F.9.
We still consider the contextual dueling bandit instance class defined in (G.11). We show that given any positive $\alpha,\eta$, for any $n\geq S\cdot\max\{16,\eta^{2}\alpha^{-2}\}$, there exists $f:\mathbb{R}\to\mathbb{R}$ such that $f$ is $\alpha$-strongly convex, $\log|\mathcal{G}|=\Theta(S)$ and
$\inf_{\widehat{\pi}\in\widehat{\Pi}(\mathcal{D})}\sup_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{f\mathrm{div}}(\widehat{\pi};\mathrm{inst})\gtrsim\frac{\eta S}{\alpha n},$ (G.16)
@@
Let ${\mathcal{S}}=[S]$, $\mathcal{A}=\{\pm 1\}$, $\rho=\mathsf{Unif}({\mathcal{S}})$ and $\pi^{\mathsf{ref}}(\cdot|s)=\mathsf{Unif}(\mathcal{A})$ for any $s\in{\mathcal{S}}$. Then the total sub-optimality of any $\pi\in\Delta(\mathcal{A}|{\mathcal{S}})$ given any reward function $r:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$ is
$\mathrm{SubOpt}_{f\mathrm{div}}(\pi;r)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{f\mathrm{div}}\big(\pi(\cdot|s);r(s,\cdot)\big).$ (G.17)
We still consider the reward function class $\mathcal{G}$ indexed by $\{\pm 1\}^{{\mathcal{S}}}$. For all $\tau\in\{\pm 1\}^{{\mathcal{S}}}$ the reward instance “shaped” by $\tau$ is
$r_{\tau}(s,a)=\frac{1}{2}+a\tau_{s}\cdot\sqrt{\frac{S}{n}},$ (G.18)
where $a\tau_{s}=\pm 1$ because $a\in\mathcal{A}=\{\pm 1\}$. We thereby refer $\tau\sim\tau^{\prime}$ to any pair in $\{\pm 1\}^{{\mathcal{S}}}$ that differs only in one coordinate. $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$, if $\tau\sim\tau^{\prime}$, then suppose $\tau_{s}=-\tau_{s}^{\prime}$, we have
$\mathrm{SubOpt}_{f\mathrm{div}}(\pi(\cdot|s);r_{\tau}(s,\cdot))+\mathrm{SubOpt}_{f\mathrm{div}}(\pi(\cdot|s);r_{\tau^{\prime}}(s,\cdot))\geq\frac{\eta S}{\alpha n},$ (G.19)
where the inequality follows from exactly the same calculation in equation (E.5) by setting $f(x)=\alpha(x-1)^{2}/2$. [^7] Recall that in this case $D_{f}={\chi^{2}}$, where $2{\chi^{2}}(\mu\|\nu)+1=\sum_{a\in\mathcal{A}}[\mu(a)]^{2}/\nu(a)$. Let $P_{r}$ be the distribution of $(s,a^{1},a^{2},y)$ for $s\sim\rho$, $a^{1},a^{2}\overset{\mathrm{i.i.d.}}{\sim}\pi^{\mathsf{ref}}(\cdot|s)$ and $y\sim\mathsf{Bern}(\sigma(r(s,a^{1})-r(s,a^{2})))$. Then we denote $\delta=\sqrt{S/n}$ and conclude that for $\tau\sim\tau^{\prime}$ with $\tau_{s}=-\tau_{s}^{\prime}$,